# First Study of the Nuclear Response to Fast Hadrons via Angular Correlations between Pions and Slow Protons in Electron–Nucleus Scattering

### v2

@@ 5 Results and discussion
Figure 2 shows the correlation function $C(\Delta\phi,\Delta Y)$ with different panels representing different $\Delta Y$ bins and the $x$ axis representing $|\Delta\phi|$ [^5] We bin in $|\Delta\phi|$ rather than $\Delta\phi$ since the observed correlation function is symmetrical with respect to $\Delta\phi=\pi$.. We note that the correlation function has a maximum, or “peak [^6] Had we shown $\Delta\phi$ from 0 to $2\pi$, as was done in Ref. PhysRevC.111.035201, this would appear as a symmetrical peak, since the correlation function is symmetrical about $\Delta\phi=\pi$. That is, the correlation function should have the same value at $\Delta\phi=179^{\circ}$ and at $\Delta\phi=181^{\circ}$.”, at $|\Delta\phi|=\pi$ (that is, where the two hadrons are azimuthally opposite one another) in every bin for every target. Further, this peak is tallest within the $1.0<\Delta Y<1.5$ bin (third column). This happens to coincide with the average rapidity separation between the $\gamma^{*}N$ system and the lab frame (about 1.2), which would be expected if the knocked-out proton rapidity distribution were peaked at that of the lab frame, and the leading pion rapidity distribution were peaked at the $\gamma^{*}N$ system. The peak at $|\Delta\phi|=\pi$ is well pronounced for projections where $\Delta Y$ is small, and more spread out in the slices where $\Delta Y$ is large.
@@
The ratios of the correlation functions for nuclei to those of deuterium are shown in the bottom row of Fig. 2 and can reach up to approximately 30 for some bins. We note that these ratios are smaller for azimuthally opposite pairs ($|\Delta\phi|\approx\pi$) than for azimuthally adjacent pairs ($|\Delta\phi|\approx 0$). This is similar to observations in dipion-production measurements CLAS:2022asf, PhysRevC.111.035201, which were attributed to absorption effects coupled with geometrical biases caused by survival selection (that is, hadrons that are absorbed in the nucleus are not detected.) Fialkowski:2007pt.
@@
We further note that, except when $\Delta Y$ is small and $|\Delta\phi|\approx\pi$, the correlation functions are larger for the nuclear targets than for deuterium (as indicated by the ratio, $R$, being greater than one). Since the correlation functions are related to the proton yield per leading pion, $N_{e^{\prime}\pi p}/N_{\pi p}$, (see Eq. 3), it follows that more protons are produced per leading-$\pi^{+}$ event in nuclei than in deuterium. This can partially be explained by the fact that there are more protons that can be knocked out of a larger nucleus than a smaller nucleus. However, this trend appears to saturate, as the values for the correlation functions for iron and lead are within about 30% of one another despite lead having more than three times as many protons as iron. In a scenario where only the protons in or near the path of the cascade could be knocked out, the number of protons that could be knocked out would scale with the nuclear radius, which can be proxied by $\sqrt[3]{A}$, rather than the atomic number, $Z$. In this case, about ($\sqrt[3]{A_{\rm Pb}}/\sqrt[3]{A_{\rm Fe}}=\sqrt[3]{208/56}\approx 1.55$ times as many protons could be knocked out from Pb than from Fe, which is closer to what we observe than if this were scaled with $Z$. Further, the cascade has a limited amount of energy, some of which causes protons to be knocked out and part of which goes into producing other particles, therefore limiting the number of protons that can be knocked out per reaction. This explanation can be tested by comparing these results to future measurements at higher energy, such as a recent experiment with the CLAS12 detector Burkert:2020akg, which used D, C, Al, Cu, Sn, and Pb targets. According to this interpretation, the saturation effect should occur at a larger value of AA for higher energies than at lower energies.
@@
We also note that in the highest $\Delta Y$ bin (rightmost column of Fig. 2), the protons and pions are very weakly correlated with one another, as indicated by the correlation function having little-to-no slope with respect to $\Delta\phi$, and that the average number of protons per pion relative to deuterium (see bottom-right panel of Fig. 2) is $\approx$2 for all three targets. Our observation that the protons are very weakly correlated with far-forward pions suggests that future measurements could, by selecting events with large or small numbers of protons, seek to enhance or decrease nuclear effects without biasing the forward pion kinematics RoblesGajardo:2022efe.
@@
We compare our results in Fig. 2 with the calculations of the BeAGLE model Chang:2022hkt (version 1.01.03), which was developed as a benchmark model for the Electron-Ion Collider Accardi:2012qut, AbdulKhalek:2021gbh. The BeAGLE event generator integrates modules from various sources. The primary interaction is managed by PYTHIA6 Sjostrand:2006za. It incorporates nuclear parton distribution functions (PDFs) from LHAPDF5 Whalley:2005nh, intranuclear scattering via DPMJet Roesler:2000he, and the geometric density of nucleons through PyQM Dupre:2011afa. Nuclear remnant de-excitations and decays are processed using FLUKA Ferrari:2005zk. We determined the correlation functions from these simulations by applying Eq. 3 to the number of simulated events within each kinematic bin. The value of $C_{0}$ used for normalizing the simulated correlation functions was determined using the deuterium simulations and is independent of the value used for data.
@@
Figure 3: Top row: RMS width $\sigma$ as a function of $\sqrt[3]{A}$ (left-most panel), and as a function of $\Delta Y$ for each target (other panels). Bottom row: same for the azimuthal broadening, $b$. The horizontal caps in the error bars represent the systematic uncertainties, while the outer bars represent the total systematic and statistical uncertainty (added in quadrature). The results are compared to the predictions from the BeAGLE Chang:2022hkt, eHIJING PhysRevD.110.034001, and GiBUU Buss:2011mx models, as well as analogous measurements of the dipion channel from Ref. PhysRevC.111.035201 (gray, open squares). The broadenings for the dipion channel were reported in Ref. PhysRevC.111.035201 with respect to deuterium, and have been re-calculated with respect to carbon in order to compare them to the results of this work.
@@
From the correlation functions, we determined the RMS widths, $\sigma$, integrated in $\Delta Y$, which we show in the top left panel of Fig. 3 as a function of $\sqrt[3]{A}$ (which proxies the radius of the nucleus). We note that $\sigma$ increases monotonically with respect to increasing nuclear size, with values of about 1.13, 1.32, 1.41, and 1.45 for D, C, Fe, and Pb, respectively. This increase is further reflected in the azimuthal broadenings, $b$, which increase as a function of AA, as shown in the lower left panel of the same figure. This trend is similar to what was observed in the dipion measurements PhysRevC.111.035201 (gray open squares in Fig. 3). We note that this increase in width with respect to nuclear size is steeper in this work than in Ref. PhysRevC.111.035201. This seems to suggest that low-energy protons undergo stronger final-state interactions than secondary pions. In the context of some models, such as the GiBUU model Buss:2011mx, when secondary pions are produced, they have a “pre-hadronic” phase when they are color neutral but are not fully formed. During this phase, the “pre-hadrons” have a weaker interaction with the nuclear material than after they have fully formed. Since protons from knockout are already fully formed, they do not have this pre-hadronic phase; therefore they interact with the nucleus as soon as they are struck by the cascade. Another possible explanation is that the protons have lower kinetic energy than pions of the same momentum, and are therefore more susceptible to energy loss and absorption effects.
@@
In the subsequent columns of this figure, we show the widths and broadenings as a function of $\Delta Y$ for each target. The widths increase monotonically as a function of $\Delta Y$ (which was also observed for the dipion channel in Ref. PhysRevC.111.035201, however we observe that the slope of this dependence is larger for the $\pi p$ channel than for the dipion channel). The broadening is consistently negative (up to $\approx$-0.8) for deuterium and positive for iron (up to $\approx 0.6$) and lead (up to $\approx 0.7$), with the exception of one bin in which the uncertainties are very large.
@@
These trends in the widths and broadenings are reproduced by the BeAGLE model, which predicts the values of $\sigma$ within about one error bar for the integrated results (upper left panel of Fig. 3). However, the agreement with the data is not as good within slices in $\Delta Y$ as in the integrated results.
@@
We also compared our results for the widths and broadenings to the predictions of the eHIJING model PhysRevD.110.034001 (version 1.0) and the GiBUU model Buss:2011mx (2021 release) in Fig. 3. We also provide a comparison between these models’ predictions for the correlation functions and the ratio, $R$, in the Supplementary Materials supplementary. In eHIJING, the hard-scattering process, along with initial- and final-state fragmentation, is simulated using PYTHIA8 bierlich2022comprehensive. The model incorporates nuclear modifications to PDFs, fragmentation, and parton showering. It also handles multiple collisions between partons and the rest of the nucleus, with a cross section proportional to the transverse-momentum-dependent gluon distribution density. Like the BeAGLE model, eHIJING was developed for the Electron-Ion Collider.
@@
The GiBUU model incorporates treatment of final-state interactions, absorption, and production mechanisms with elastic and inelastic channels. The GiBUU model describes the single-hadron and di-hadron data reasonably well from CLAS CLAS:2021jhm, CLAS:2022asf, PhysRevC.111.035201 and HERMES HERMES:2003icw, Airapetian:2007vu, Airapetian:2011jp. We modified the default values of several parameters related to parton distributions and fragmentation in free nucleons, based on a tuning study that used an independent proton-target dataset from Ref. CLAS:2022sqt, and which was further validated with dipion correlations in Ref. PhysRevC.111.035201. This tuning does not modify the nuclear effects.
@@ 6 Summary and conclusions
The azimuthal‑peak width grows systematically with nuclear size, and the proton yield per pion also rises but it appears to saturate for the heaviest targets. These patterns are consistent with a picture in which hadron absorption and interactions in the hadronic cascade widens the correlation and the larger pool of protons in heavier nuclei enhances the yield.
Compared with earlier dipion measurements PhysRevC.111.035201, the $\pi p$ channel shows a markedly stronger dependence on both nuclear size and rapidity separation. The trends are qualitatively reproduced by the state-of-the-art eAeA event-generator models BeAGLE, eHIJING, and GiBUU, but quantitative differences remain, highlighting opportunities to refine the treatment of nuclear fragmentation.
Our results establish a benchmark for forthcoming di-hadron and slow-nucleon production studies in DIS, as well as for the development of corresponding event-generator models for experiments at Jefferson Lab Burkert:2018nvj, Burkert:2020akg, Arrington:2021alx and at future Electron–Ion Colliders Accardi:2012qut, Anderle:2021wcy.
@@ 7 Code and Data Availability
We provide in Ref. zenodoRivet the code we used to extract the $(\Delta\phi,\Delta Y)$ histograms from the output files of event generators (in order to produce the correlation functions) in the form of a module in the Rivet framework 10.21468/SciPostPhys.8.2.026.
The values of the correlation functions, the ratio $R$, RMS widths, and broadenings reported in this work can be found in the Supplementary Materials supplementary.
@@ 8 Acknowledgements
The authors acknowledge the staff of the Accelerator and Physics Divisions at the Thomas Jefferson National Accelerator Facility who made this experiment possible. We thank Kai Gallmeister for help in setting up the GiBUU event generator and Weiyao Ke for help in setting up the eHIJING event generator. This work is supported by the Chilean Agencia Nacional de Investigacion y Desarollo (ANID), FONDECYT Grants No. 1221827 and No. 1240904, ANID PIA/APOYO AFB230003 and by the ANID-Millennium Science Initiative Program–ICN2019_044, by the U.S. Department of Energy, the Italian Instituto Nazionale di Fisica Nucleare, the French Centre National de la Recherche Scientifique, the French Commissariat à l’Energie Atomique, the United Kingdom Science and Technology Facilities Council (STFC), the Scottish Universities Physics Alliance (SUPA), the National Research Foundation of Korea, the National Science Foundation (NSF), the HelmholtzForschungsakademie Hessen für FAIR (HFHF), the Ministry of Science and Higher Education of the Russian Federation, and the Office of Research and Economic Development at Mississippi State University. This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 804480). The Southeastern Universities Research Association operates the Thomas Jefferson National Accelerator Facility for the United States Department of Energy under Contract No. DE-AC05-06OR23177.

### v3

@@ 5 Results and discussion
Figure 2 shows the correlation function $C(\Delta\phi,\Delta Y)$ with different panels representing different $\Delta Y$ bins and the $x$ axis representing $|\Delta\phi|$ [^5] We bin in $|\Delta\phi|$ rather than $\Delta\phi$ since the observed correlation function is symmetrical with respect to $\Delta\phi=\pi$.. We note that the correlation function has a maximum, or “peak [^6] Had we shown $\Delta\phi$ from 0 to $2\pi$, as was done in Ref. Paul and others [2025], this would appear as a symmetrical peak, since the correlation function is symmetrical about $\Delta\phi=\pi$. That is, the correlation function should have the same value at $\Delta\phi=179^{\circ}$ and at $\Delta\phi=181^{\circ}$.”, at $|\Delta\phi|=\pi$ (that is, where the two hadrons are azimuthally opposite one another) in every bin for every target. Further, this peak is tallest within the $1.0<\Delta Y<1.5$ bin (third column). This happens to coincide with the average rapidity separation between the $\gamma^{*}N$ system and the lab frame (about 1.2), which would be expected if the knocked-out proton rapidity distribution were peaked at that of the lab frame (as would be the case if their initial momentum distribution is isotropic), and the leading pion rapidity distribution were peaked at the $\gamma^{*}N$ system from which it originated. This interpretation could be tested by seeing if this observation continues to hold true in future measurements, e.g in experiments with the CLAS12 detector Burkert and others [2020], where a larger range in $\nu$ and $Q^{2}$ allows a larger range of values for the rapidity of the $\gamma^{*}N$ system. The peak at $|\Delta\phi|=\pi$ is well pronounced for projections where $\Delta Y$ is small, and more spread out in the slices where $\Delta Y$ is large.
@@
The ratios of the correlation functions for nuclei to those of deuterium are shown in the bottom row of Fig. 2. The values of these ratios range from near unity to about 30, depending on the target and kinematic bin. We note that these ratios are smaller for azimuthally opposite pairs ($|\Delta\phi|\approx\pi$) than for azimuthally adjacent pairs ($|\Delta\phi|\approx 0$). This is similar to observations in dipion-production measurements Paul and others [2022, 2025], which were attributed to absorption effects coupled with geometrical biases caused by survival selection (that is, hadrons that are absorbed in the nucleus are not detected.) Fialkowski and Wit [2007].
@@
We further note that, except when $\Delta Y$ is small and $|\Delta\phi|\approx\pi$, the correlation functions are larger for the nuclear targets than for deuterium (as indicated by the ratio, $R$, being greater than one). Since the correlation functions are related to the proton yield per leading pion, $N_{e^{\prime}\pi p}/N_{e^{\prime}\pi}$, (see Eq. 3), it follows that more protons are produced per leading-$\pi^{+}$ event in nuclei than in deuterium. This can partially be explained by the fact that there are more protons that can be knocked out of a larger nucleus than a smaller nucleus. However, this trend appears to saturate, as the values for the correlation functions for iron and lead are within about 30% of one another despite lead having more than three times as many protons as iron. In a scenario where only the protons in or near the path of the cascade could be knocked out, the number of protons that could be knocked out would scale with the nuclear radius, which can be proxied by $\sqrt[3]{A}$, rather than the atomic number, $Z$. In this case, about $\sqrt[3]{A_{\rm Pb}}/\sqrt[3]{A_{\rm Fe}}=\sqrt[3]{208/56}\approx 1.55$ times as many protons could be knocked out from Pb than from Fe, which is closer to what we observe than if this were scaled with $Z$. Further, the cascade has a limited amount of energy, some of which causes protons to be knocked out and part of which goes into producing other particles, therefore limiting the number of protons that can be knocked out per reaction. This explanation can be tested by comparing these results to future measurements at higher energy, such as a recent experiment with the CLAS12 detector Burkert and others [2020], which used D, C, Al, Cu, Sn, and Pb targets. According to this interpretation, the saturation effect should occur at a larger value of $A$ for higher energies than at lower energies.
@@
We also note that in the highest $\Delta Y$ bin (rightmost column of Fig. 2), the protons and pions are very weakly correlated with one another, as indicated by the correlation function having little-to-no slope with respect to $\Delta\phi$, and that the average number of protons per pion relative to deuterium (see bottom-right panel of Fig. 2) is $\approx$2 for all three targets. Our observation that the protons are very weakly correlated with far-forward pions suggests that future measurements could, by selecting events with large or small numbers of protons, seek to enhance or decrease nuclear effects without biasing the forward pion kinematics Robles Gajardo et al. [2022].
@@
We compare our results in Fig. 2 with the calculations of the BeAGLE model Chang et al. [2022] (version 1.01.03), which was developed as a benchmark model for the Electron-Ion Collider Accardi and others [2016], Abdul Khalek and others [2022]. The BeAGLE event generator integrates modules from various sources. The primary interaction is managed by PYTHIA6 Sjostrand et al. [2006]. It incorporates nuclear parton distribution functions (PDFs) from LHAPDF5 Whalley et al. [2005], intranuclear scattering via DPMJet Roesler et al. [2000], and the geometric density of nucleons through PyQM Dupré [2011]. Nuclear remnant de-excitations and decays are processed using FLUKA Ferrari et al. [2005]. We determined the correlation functions from these simulations by applying Eq. 3 to the number of simulated events within each kinematic bin. The value of $C_{0}$ used for normalizing the simulated correlation functions was determined using the deuterium simulations and is independent of the value used for data. However, they are close to one another: $C_{0}^{\rm data}\approx 7.5$ and $C_{0}^{\rm BeAGLE}\approx 8.1$, indicating that the prediction for the average number of ejected protons per leading-pion event in deuterium is within 10% of that of the data.
@@
Figure 3: Top row: RMS width $\sigma$ as a function of $\sqrt[3]{A}$ (left-most panel), and as a function of $\Delta Y$ for each target (other panels). Bottom row: same for the azimuthal broadening, $b$. The horizontal caps in the error bars represent the systematic uncertainties, while the outer bars represent the total systematic and statistical uncertainty (added in quadrature). The results are compared to the predictions from the BeAGLE Chang et al. [2022], eHIJING Ke et al. [2024], and GiBUU Buss et al. [2012] models, as well as analogous measurements of the dipion channel from Ref. Paul and others [2025] (gray, open squares). The broadenings for the dipion channel were reported in Ref. Paul and others [2025] with respect to deuterium, and have been re-calculated with respect to carbon in order to compare them to the results of this work.
@@
From the correlation functions, we determined the RMS widths, $\sigma$, integrated in $\Delta Y$, which we show in the top left panel of Fig. 3 as a function of $\sqrt[3]{A}$ (which proxies the radius of the nucleus). We note that $\sigma$ increases monotonically with respect to increasing nuclear size, with values of about 1.13, 1.32, 1.41, and 1.45 for D, C, Fe, and Pb, respectively. This increase is further reflected in the azimuthal broadenings, $b$, which increase as a function of $A$, as shown in the lower left panel of the same figure. This trend is similar to what was observed in the dipion measurements Paul and others [2025] (gray open squares in Fig. 3). We note that this increase in width with respect to nuclear size is steeper in this work than in Ref. Paul and others [2025]. This seems to suggest that low-energy protons undergo stronger final-state interactions than secondary pions. In the context of some models, such as the GiBUU model Buss et al. [2012], when secondary pions are produced, they have a “pre-hadronic” phase when they are color neutral but are not fully formed. During this phase, the “pre-hadrons” have a weaker interaction with the nuclear material than after they have fully formed. Since protons from knockout are already fully formed, they do not have this pre-hadronic phase; therefore they interact with the nucleus as soon as they are struck by the cascade. Another possible explanation is that the protons have lower kinetic energy than pions of the same momentum, and are therefore more susceptible to energy loss and absorption effects.
@@
In the subsequent columns of this figure, we show the widths and broadenings as a function of $\Delta Y$ for each target. The widths increase monotonically as a function of $\Delta Y$ (which was also observed for the dipion channel in Ref. Paul and others [2025], however we observe that the slope of this dependence is larger for the $\pi p$ channel than for the dipion channel). The broadening is consistently negative (up to $\approx$-0.8) for deuterium and positive for iron (up to $\approx 0.6$) and lead (up to $\approx 0.7$), with the exception of one bin in which the uncertainties are very large.
@@
These trends in the widths and broadenings are reproduced by the BeAGLE model, which predicts the values of $\sigma$ within about one error bar for the integrated results (upper left panel of Fig. 3). However, the agreement with the data is not as good within slices in $\Delta Y$ as in the integrated results. This appears to be related to the discrepancy mentioned above, where the predicted correlation functions appear to be shifted in $\Delta Y$ relative to the data.
@@
We also compared our results for the widths and broadenings to the predictions of the eHIJING model Ke et al. [2024] (version 1.0) and the GiBUU model Buss et al. [2012] (2021 release) in Fig. 3. We also provide a comparison between these models’ predictions for the correlation functions and the ratio, $R$, in the Supplementary Materials 1. In eHIJING, the hard-scattering process, along with initial- and final-state fragmentation, is simulated using PYTHIA8 Bierlich et al. [2022]. The model incorporates nuclear modifications to PDFs, fragmentation, and parton showering. It also handles multiple collisions between partons and the rest of the nucleus, with a cross section proportional to the transverse-momentum-dependent gluon distribution density. Like the BeAGLE model, eHIJING was developed for the Electron-Ion Collider.
@@
The GiBUU model incorporates treatment of final-state interactions, absorption, and production mechanisms with elastic and inelastic channels. The GiBUU model describes the single-hadron and di-hadron data reasonably well from CLAS Moran and others [2022], Paul and others [2022, 2025] and HERMES Airapetian and others [2003, 2007, 2011]. We modified the default values of several parameters related to parton distributions and fragmentation in free nucleons, based on a tuning study that used an independent proton-target dataset from Ref. Avakian and others [2023], and which was further validated with dipion correlations in Ref. Paul and others [2025]. This tuning does not modify the nuclear effects.
@@ 6 Summary and conclusions
The azimuthal‑peak width grows systematically with nuclear size, and the proton yield per pion also rises but it appears to saturate for the heaviest targets. These patterns are consistent with a picture in which hadron absorption and interactions in the hadronic cascade widen the correlation and the larger pool of protons in heavier nuclei enhances the yield.
Compared with earlier dipion measurements Paul and others [2025], the $\pi p$ channel shows a markedly stronger dependence on both nuclear size and rapidity separation. The trends are qualitatively reproduced by the state-of-the-art $eA$ event-generator models BeAGLE, eHIJING, and GiBUU, but quantitative differences remain, highlighting opportunities to refine the treatment of nuclear fragmentation.
Our results establish a benchmark for forthcoming di-hadron and slow-nucleon production studies in DIS, as well as for the development of corresponding event-generator models for experiments at Jefferson Lab Burkert [2018], Burkert and others [2020], Arrington and others [2022] and at future Electron–Ion Colliders Accardi and others [2016], Anderle and others [2021].
@@ 7 Code and Data Availability
We provide in Ref. Paul the code we used to extract the $(\Delta\phi,\Delta Y)$ histograms from the output files of event generators (in order to produce the correlation functions) in the form of a module in the Rivet framework Bierlich et al. [2020].
The values of the correlation functions, the ratio $R$, RMS widths, and broadenings reported in this work can be found in the Supplementary Materials 1.
@@ 8 Acknowledgements
The authors acknowledge the staff of the Accelerator and Physics Divisions at the Thomas Jefferson National Accelerator Facility who made this experiment possible. We thank Kai Gallmeister for help in setting up the GiBUU event generator and Weiyao Ke for help in setting up the eHIJING event generator. This work is supported by the Chilean Agencia Nacional de Investigación y Desarrollo (ANID), FONDECYT Grants No. 1221827 and No. 1240904, ANID PIA/APOYO AFB230003 and by the ANID-Millennium Science Initiative Program – ICN2019_044, by the U.S. Department of Energy, the Italian Instituto Nazionale di Fisica Nucleare, the French Centre National de la Recherche Scientifique, the French Commissariat à l’Energie Atomique, the United Kingdom Science and Technology Facilities Council (STFC), the Scottish Universities Physics Alliance (SUPA), the National Research Foundation of Korea, the National Science Foundation (NSF), the HelmholtzForschungsakademie Hessen für FAIR (HFHF), the Ministry of Science and Higher Education of the Russian Federation, and the Office of Research and Economic Development at Mississippi State University. This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 804480). The Southeastern Universities Research Association operates the Thomas Jefferson National Accelerator Facility for the United States Department of Energy under Contract No. DE-AC05-06OR23177.